# Casimir-Lifshitz interaction between bodies integrated in a micro/nanoelectromechanical quantum damped oscillator

### v1

@@ Abstract
Casimir-Lifshitz forces in circuits.Assuming that dissipation in electrical circuits is described within the Zwanzig-Caldeira-Leggett model [65, 66, 67], we apply the above results to the two simplest RLC circuits: one in series and the other in parallel, as depicted in Fig. 1.
@@
In the RLC circuit connected in series, the charge $Q$ can be taken as the system’s generalized coordinate [68]. The corresponding canonical momentum is the current-induced magnetic flux $\Phi=L\dot{Q}$. The equation of motion for the circuit, $L\ddot{Q}+R\dot{Q}+\frac{Q}{C}={\cal E}_{ext}(t)$, with the external voltage ${\cal E}_{ext}(t)$, has the same form as that of the damped harmonic oscillator. In this analogy, $M\rightarrow L$, $\gamma\rightarrow\frac{R}{L}$ and $\Omega\rightarrow\frac{1}{\sqrt{LC}}$. Note that $\gamma$ here represents the frequency-independent Ohmic damping.
@@
If the capacitance $C$ depends on the distance $d$, while $R$ and $L$ do not, then only the oscillator frequency $\Omega$ depends on $d$, whereas $\gamma$ and $M$ do not. Therefore, the contribution $f_{RLC}$ of electromagnetic fluctuations in the series circuit to the Casimir-Lifshitz force is determined by equation (2), with parameters $\Omega_{LC}=\frac{1}{\sqrt{LC}}$ and $\gamma=\frac{R}{L}$. Alternatively, if $C$ is the only circuit parameter that depends on a quantity $\varkappa$, while some other parameters (possibly including $C$) depend on $d$, then the finite force difference is $\tilde{f}_{RLC_{1}}-\tilde{f}_{RLC_{2}}$, where $\tilde{f}_{RLC}$ is defined in (4) with $\Omega_{LC_{1,2}}=\frac{1}{\sqrt{LC_{1,2}}}$, $\gamma=\frac{R}{L}$ and $C_{1,2}=C(\varkappa_{1,2})$.
@@
In the RLC circuit connected in parallel, the generalized coordinate is $\Phi=\int_{-\infty}^{t}V(t^{\prime})dt^{\prime}$, where $\Phi$ is the magnetic flux. The capacitor charge $Q_{C}=C\dot{\Phi}=CV$ serves as the canonical momentum in this case [65, 67, 69]. The corresponding equation of motion, $C\ddot{\Phi}(t)+\dfrac{1}{R}\dot{\Phi}(t)+\dfrac{1}{L}\Phi(t)=J_{\text{ext}}(t)$, coincides with that for a damped harmonic oscillator, where $M\rightarrow C$, $\gamma\rightarrow\frac{1}{RC}$ and $\Omega\rightarrow\frac{1}{\sqrt{LC}}$. Here, as before, $\gamma$ is frequency-independent.
@@
If the inductance $L$ depends on distance $d$, while $R$ and $C$ do not, then only the oscillator frequency $\Omega$ in the parallel RLC circuit depends on $d$, whereas $\gamma$ and $M$ remain unchanged. Therefore, the contribution $f_{RLC}$ of the circuit to the Casimir-Lifshitz force in the parallel configuration is determined by (2), with parameters $\Omega_{LC}=\tfrac{1}{\sqrt{LC}}$ and $\gamma=\tfrac{1}{RC}$. Alternatively, if $L$ is the only circuit parameter that depends on a quantity $\varkappa$, while a few circuit parameters (possibly, including $L$) depend on $d$, then the force difference is $f_{RL_{1}C}-f_{RL_{2}C}=\tilde{f}_{RL_{1}C}-\tilde{f}_{RL_{2}C}$, where $\tilde{f}_{RLC}$ is defined in (4), $\Omega_{LC_{1,2}}=\tfrac{1}{\sqrt{L_{1,2}C}}$, $\gamma=\tfrac{1}{RC}$ and $L_{1,2}=L(\varkappa_{1,2})$.
@@
We estimate the Casimir-Lifshitz force component $f_{RLC}$ for the RLC loop in series involving the distance-dependent capacitance, and examine further the ratio $r=\frac{f_{RLC}}{f_{Cas}}$, which compares this force component to the main Casimir-Lifshitz force $f_{Cas}$ between the same bodies under similar conditions. Consider first the planar capacitor, for which the resonant frequency of the circuit is $\Omega^{pc}_{LC}=\sqrt{\frac{d}{\varepsilon_{0}\varepsilon LS}}$, where $S$ and $d$ denote the contact area of the conducting plates and the distance between them, respectively; $\varepsilon_{0}$ is the vacuum permittivity. Focusing on the Casimir problem [1], we put $\varepsilon=1$ for the dielectric permittivity of the interlayer between the plates.
@@
The force at high temperatures described by the first term in (2), $f_{T}=-\frac{T}{\Omega}\frac{\partial\Omega}{\partial d}$, takes the form $f^{pc}_{RLC,T}=-\frac{T}{2d}$ in this case. At low temperatures, one obtains from (3) for the interaction force at weak ($R^{2}\ll\frac{4Ld}{\varepsilon S}$) and strong ($R^{2}\gg\frac{4Ld}{\varepsilon S}$) dissipation, respectively:
@@
Given the applicability conditions of the lumped element approach, the strength of dissipation in (8) is confined by the requirement $i\omega_{1}\approx\frac{R}{L}\ll\min(\omega_{c},c/r_{0})$, where $r_{0}$ is the typical size of a circuit element.
Furthermore, the conventional Casimir force, $f_{Cas}=-\frac{\pi^{2}\hbar cS}{240d^{4}}$, is the main fluctuation force component between metal plates in the strong retardation regime at low temperatures. At high temperatures, $T\gg\hbar c/d$, the force becomes $f_{Cas,T}=-\frac{\zeta(3)TS}{8\pi d^{3}}$. This expression follows from the Lifshitz theory with the Drude model  [^2] As was noted long ago (see p.263 in Ref. [86]), this is half of what follows from the plasma model, since the latter fails to account for the complete suppression of the TM polarization contribution at high temperatures. This fact was later rediscovered and drew considerable attention due to a reported discrepancy between precision measurements and predictions from Lifshitz theory [90, 91, 77, 26, 92].
One finds from here the relative weights $r^{pc}_{0}$ and $r^{pc}_{T}$ of the circuit-induced Casimir-Lifshitz force between the plates of the planar capacitor in the dissipationless limit, $\hbar\gamma\ll\hbar\Omega,T$, at low and high temperatures:
@@
As seen in (9) and (10), the relative weight involves two small parameters at low temperature range and only one at high. The quantities $r^{pc}_{0}$ and $r_{\phantom{}{}_{T}}^{pc}$, are both governed by the small geometric factor $\frac{d^{2}}{S}$. It reflects the different dependence of the forces on the separation between the bodies and their characteristic sizes. Its smallness justifies neglecting edge effects at the boundaries of the plates.
Another small parameter, $\frac{\Omega^{pc}_{LC}(d)}{c/d}$, that appears in $r_{0}^{pc}$ is the ratio of the oscillator frequency $\Omega^{pc}_{LC}$ and the characteristic frequency $c/d$ of quantum fluctuations forming the Casimir force. This ratio is typically small throughout the entire applicability domain of the Casimir result $f_{Cas}$. After switching over from low to high temperatures, both characteristic frequencies, $\Omega^{pc}_{LC}$ and $d/c$, present in (9) are effectively replaced by $T$, which cancels out the frequencies in the relative weight, leading to (10), up to a numerical factor.
As the ratio $\frac{d^{2}}{S}$ decreases from $0.04$ to $2.5\!\cdot\!10^{-3}$, the relative weight $r^{pc}_{\phantom{}{}_{T}}$ in (10) drops from $0.42$ to $0.03$. The corresponding measurement accuracy requirements indicate that the force component under consideration can generally be detected at high temperatures.
For estimating the relative weights for a metal sphere and a flat metal plate often used in experimental setups [12, 13, 71, 30, 14, 14, 72, 73, 74, 75, 76, 18, 77, 22, 41, 23, 25, 47], the capacitance could be calculated numerically [78, 79, 80, 81]. However, we find it sufficient for our purposes to rely on a simple interpolating expression, $C^{sp-p}(d,R_{sp})=4\pi\varepsilon_{0}R_{sp}\left[1+\frac{1}{2}\log\left(1+\frac{R_{sp}}{d}\right)\right]$, which provides a good approximation for $d\lesssim R_{sp}$ [80]. Here, $R_{sp}$ is the radius of the sphere, and $d$ is the minimum sphere-plate distance. In the dissipationless limit and at low temperatures, $T\ll\hbar\Omega^{sp-p}_{LC}$, the circuit-induced fluctuation force between the sphere and the plate, associated with the corresponding $\Omega^{sp-p}_{LC}=\frac{1}{\sqrt{LC^{sp-p}}}$, is
@@
The factor $\Omega^{sp-p}_{LC}/(c/d)$ that enters (13) is typically small within the applicability domain of the Casimir result. However, this factor is canceled out in (14). The geometric parameter $\frac{d}{R_{sp}}$ can vary here within comparatively wide limits. The relative weight $r^{sp-p}_{T}$ decreases from $0.5$ to $0.02$ as $\frac{d}{R_{sp}}$ decreases from $0.75$ to $0.035$. This shows a possibility to detect the circuit-induced force $f^{sp-p}_{LC}$ under optimal conditions. The presence of multiple distance-dependent eigenmodes could enhance the effect.
Discussion and conclusions.Components of the Casimir-Lifshitz forces induced in electrical circuits were previously examined in Ref. [83]. That paper addresses a number of important aspects of the issue under discussion and presents a series of valuable and useful findings. However, the main results and statements in Ref. [83] concerning the circuit-induced component of the Casimir-Lifshitz force are in contradiction to those obtained here. This relates both to the formulas for the forces in the circuits and to the more general results and conclusions.
One such conclusion in Ref. [83] is that, in the dissipationless limit at zero temperature, the interaction potential differs from the zero-point energy. The other conclusion is that the divergence of the interaction potential, as described within the lumped element approach, also entails the divergence of the interaction force, even when it is only the resonant frequency that depends on distance. Specifically, the expression for interaction force in the series RLC loop, as presented in Ref. [83], is divergent in the case of a distance-dependent $C$ at constant $R$ and $L$. This contradicts our fomulas (7) and (8), as well as the more general results (2) and (3).
The main reason for the contradictions is that the force component between the capacitor plates, as considered in Ref. [83], is attributed solely to charge fluctuations on the plates themselves, without taking into account other contributions. However, this only represents part of the total Casimir-Lifshitz force induced by the circuit, since changing the plate separation alters the impedance of the entire circuit. Consequently, the circuit’s free energy dependence on distance is determined by fluctuations in all lumped circuit elements. In this regard, it is worth recalling that the fluctuation-induced interaction between oscillators or atoms originates from both potential and kinetic fluctuational energy terms [84, 85].
Before the technological breakthrough in the field, the fluctuation-induced forces in the series RLC circuit were considered a purely methodological issue [86],  [^3] See also the footnote on p.407 in Ref. [8]., primarily in connection with Ref. [88]. While the fluctuation-induced free energy in the series RLC circuit diverges, it was shown that the difference between free energies, taken at different capacitance values, is finite. A finite additional component of the fluctuation force that was identified in Ref. [86] for the circuit, coincides with (7), (8) up to a minor typo  [^4] In the limit of large dissipation, described by equation (5.422) on page 248 in Ref. [86], the resistance $R$ under the logarithm sign must be squared.. Those early results were obtained by simply expressing the fluctuation energy in the RLC circuit as the average of $\frac{L\dot{Q}^{2}}{2}+\frac{Q^{2}}{2C}$.

### v2

@@ III Casimir-Lifshitz forces in circuits
Assuming that dissipation in electrical circuits can be described within the Zwanzig-Caldeira-Leggett model [66, 67, 68], we apply the above results for the quantum damped oscillator to the damped eigenmodes of electrical circuits. To this end, we identify the resonant frequency and damping function for the particular circuit mode, and then substitute these in (1)–(4). In the general theory of van der Waals forces, the expansion over the damped eigenmodes is developed in detail [8]. Two particular RLC circuits, one in series and the other in parallel, are considered here as the simplest examples depicted in Fig. 1.
@@
In series RLC circuit, the current $J$ and the corresponding charge $Q$ ($J=\dot{Q}$) are common to all lumped elements. Therefore, it is convenient to take $Q$ as a generalised coordinate of the system [69]. The corresponding canonical momentum that enters the Hamiltonian of the full conservative system along with the coordinates and momenta of the environment, is the current-induced magnetic flux $\Phi=L\dot{Q}$. The effective equation of motion for the circuit, obtained after eliminating the environmental variables, is $L\ddot{Q}+R\dot{Q}+\tfrac{Q}{C}={\cal E}_{ext}(t)$, where ${\cal E}_{ext}(t)$ is the external voltage. This equation describes the circuit dynamics and has the same form as that of a damped harmonic oscillator. In this analogy, $M\rightarrow L$, $\gamma\rightarrow\tfrac{R}{L}$ and $\Omega\rightarrow\tfrac{1}{\sqrt{LC}}$. Note that $\gamma$ here represents the frequency-independent Ohmic damping.
@@
If the capacitance $C$ depends on the distance $d$, while $R$ and $L$ do not, then only the oscillator frequency $\Omega$ depends on $d$, whereas $\gamma$ and $M$ do not. The contribution $f_{RLC}$ of electromagnetic fluctuations in the series circuit to the Casimir-Lifshitz force is then given by equation (2), with parameters $\Omega_{LC}=\tfrac{1}{\sqrt{LC}}$ and $\gamma=\tfrac{R}{L}$.
Alternatively, assume that the interacting bodies make a noticeable contribution to the circuit inductance, which therefore becomes dependent on the distance $d$ between them, while the capacitance is mainly determined by other circuit components. In this case, the capacitance can be varied at a fixed interaction, in particular, at fixed $L$. The finite force difference is then $\tilde{f}_{RLC_{1}}-\tilde{f}_{RLC_{2}}$, where $\tilde{f}_{RLC}$ is defined in (4) with $\Omega_{LC_{1,2}}=\tfrac{1}{\sqrt{LC_{1,2}}}$, $\gamma=\tfrac{R}{L}$ and $C_{1,2}=C(\varkappa_{1,2})$ [65]. The parameter $\varkappa$ may, for example, correspond to two different distances between the capacitor plates (unrelated now to the interbody distance) or to two different values of the dielectric permittivity of the spacer.
@@
In parallel RLC circuit, the applied voltage and the magnetic flux $\Phi=\int_{-\infty}^{t}V(t^{\prime})dt^{\prime}$ are common to all lumped elements. Therefore, it is convenient to choose $\Phi$ as the generalized coordinate of the system. The capacitor charge $Q_{C}=C\dot{\Phi}=CV$ serves as the canonical momentum in this case [66, 68, 70]. The effective equation of motion for the circuit, $C\ddot{\Phi}(t)+\tfrac{1}{R}\dot{\Phi}(t)+\tfrac{1}{L}\Phi(t)=J_{\text{ext}}(t)$, coincides with that of a damped harmonic oscillator, where $M\rightarrow C$, $\gamma\rightarrow\tfrac{1}{RC}$ and $\Omega\rightarrow\tfrac{1}{\sqrt{LC}}$. Here, as before, $\gamma$ is frequency-independent.
@@
If the inductance $L$ depends on distance $d$, while $R$ and $C$ do not, then only the oscillator frequency $\Omega$ in the parallel RLC circuit depends on $d$, whereas $\gamma$ and $M$ remain unchanged. Therefore, the contribution $f_{RLC}$ of the circuit to the Casimir-Lifshitz force in the parallel configuration is determined by (2), with parameters $\Omega_{LC}=\tfrac{1}{\sqrt{LC}}$ and $\gamma=\tfrac{1}{RC}$.
Alternatively, assume that the interacting bodies make a noticeable contribution to the circuit capacitance, which thus becomes dependent on the distance $d$ between them, while the inductance is mainly determined by other circuit components. In this case, the $L$ can be varied independently at fixed $d$ and $C$. The finite force difference is then $\tilde{f}_{RL_{1}C}-\tilde{f}_{RL_{2}C}$, where $\tilde{f}_{RLC}$ is defined in (4) with $\Omega_{L_{1,2}C}=\tfrac{1}{\sqrt{L_{1,2}C}}$, $\gamma=\tfrac{1}{RC}$, and $L_{1,2}=L(\varkappa_{1,2})$, corresponding, for example, to different spacing between the turns or to different numbers of the turns [65].
@@
We now estimate the Casimir-Lifshitz force component $f_{RLC}$ for a series RLC loop with distance-dependent capacitance and examine the ratio $r=\tfrac{f_{RLC}}{f_{Cas}}$, which compares this component to the main Casimir-Lifshitz force $f_{Cas}$ between the same bodies under similar conditions.
Consider first a planar capacitor, for which the resonant frequency is $\Omega^{pc}_{LC}=\sqrt{\tfrac{d}{\varepsilon_{0}\varepsilon LS}}$, where $S$ and $d$ denote the contact area of the conducting plates and the distance between them, respectively; $\varepsilon_{0}$ is the vacuum permittivity. Focusing on the Casimir problem [1], we set $\varepsilon=1$ for the permittivity of the interlayer between the plates.
@@
The force at high temperatures described by the first term in (2), $f_{T}=-\tfrac{T}{\Omega}\tfrac{\partial\Omega}{\partial d}$, becomes $f^{pc}_{RLC,T}=-\tfrac{T}{2d}$ in this case. At low temperatures, one obtains from (3) at weak ($R^{2}\ll\tfrac{4Ld}{\varepsilon S}$) and strong ($R^{2}\gg\tfrac{4Ld}{\varepsilon S}$) dissipation, respectively:
@@
Given the applicability conditions of the lumped element approach, the strength of dissipation in (8) is confined by the requirement $i\omega_{1}\approx\tfrac{R}{L}\ll\min(\omega_{c},c/r_{0})$, where $r_{0}$ is the typical size of a circuit element.
The conventional Casimir force, $f_{Cas}=-\tfrac{\pi^{2}\hbar cS}{240d^{4}}$, is the main fluctuation force component between metal plates in the strong retardation regime at low temperatures. At high temperatures, $T\gg\hbar c/d$, the force becomes $f_{Cas,T}=-\tfrac{\zeta(3)TS}{8\pi d^{3}}$. This expression follows from the Lifshitz theory with the Drude model  [^3] As was noted long ago (see p.263 in Ref. [86]), this is half of what follows from the plasma model, since the latter fails to account for the complete suppression of the TM polarization contribution at high temperatures. This fact was later rediscovered and drew considerable attention due to a reported discrepancy between precision measurements and predictions from Lifshitz theory [94, 95, 78, 26, 96].
From these expressions we obtain the relative weights $r^{pc}_{0}$ and $r^{pc}_{T}$ of the circuit-induced Casimir-Lifshitz force between the plates of the planar capacitor in the dissipationless limit, $\hbar\gamma\ll\hbar\Omega,T$, at low and high temperatures:
@@
As seen in (9) and (10), the relative weight involves two small parameters at low temperature range, and only one at high. The quantities $r^{pc}_{0}$ and $r_{\phantom{}{}_{T}}^{pc}$, are both governed by the small geometric factor $\tfrac{d^{2}}{S}$. It reflects the different dependence of the forces on the separation between the bodies and their characteristic sizes. Its smallness justifies neglecting edge effects at the boundaries of the plates.
Another small parameter, $\tfrac{\Omega^{pc}_{LC}(d)}{c/d}$, that appears in $r_{0}^{pc}$ is the ratio of the oscillator frequency $\Omega^{pc}_{LC}$ and the characteristic frequency $c/d$ of quantum fluctuations forming the Casimir force. This ratio is typically small throughout the entire applicability domain of the Casimir result $f_{Cas}$. After switching over from low to high temperatures, both characteristic frequencies, $\Omega^{pc}_{LC}$ and $d/c$, present in (9) are effectively replaced by $T$, which cancels out the frequencies in the relative weight, leading to (10), up to a numerical factor.
As $\tfrac{d^{2}}{S}$ decreases from $0.04$ to $2.5\!\cdot\!10^{-3}$, the relative weight $r^{pc}_{\phantom{}{}_{T}}$ in (10) drops from $0.42$ to $0.03$. The corresponding measurement accuracy requirements indicate that the force component under consideration can generally be detected at high temperatures.
For estimating the relative weights for a metal sphere and a flat metal plate, often used in experimental setups [12, 13, 72, 30, 14, 14, 73, 74, 75, 76, 77, 18, 78, 22, 41, 23, 25, 47], the capacitance could be computed numerically [79, 80, 81, 82]. However, for our purposes a simple interpolation $C^{sp-p}(d,R_{sp})=4\pi\varepsilon_{0}R_{sp}\left[1+\tfrac{1}{2}\log\left(1+\tfrac{R_{sp}}{d}\right)\right]$ provides an adequate approximation for $d\lesssim R_{sp}$ [81]. Here, $R_{sp}$ is the sphere radius and $d$ is the minimum sphere-plate separation. In the dissipationless limit and at low temperatures, $T\ll\hbar\Omega^{sp-p}_{LC}$, the circuit-induced fluctuation force between the sphere and the plate, associated with the corresponding resonant frequency $\Omega^{sp-p}_{LC}=\tfrac{1}{\sqrt{LC^{sp-p}}}$, is
@@
The factor $\Omega^{sp-p}_{LC}/(c/d)$ that enters (13) is typically small within the applicability domain of the Casimir result. However, this factor is canceled out in (14). The geometric parameter $\tfrac{d}{R_{sp}}$ can vary here within comparatively wide limits. The relative weight $r^{sp-p}_{T}$ decreases from $0.5$ to $0.02$ as $\tfrac{d}{R_{sp}}$ decreases from $0.75$ to $0.035$. This shows a possibility to detect the circuit-induced force $f^{sp-p}_{LC}$ under optimal conditions. The presence of multiple distance-dependent eigenmodes could enhance the effect.
IV Discussion
The lumped-element form of the equations of motion for electrical circuits is known to follow from a microscopic approach under certain conditions in a relatively low-frequency range. In this case, the circuit parameters $C$, $L$ and $R$ can be regarded as phenomenological constants. To determine their specific form and values in a given system constitutes an independent task.
As damped oscillators and dissipative circuits are open, nonconservative systems, a phenomenological description of their equilibrium energies and fluctuation-induced forces must, in general, be additionally validated by a microscopic theory. The Zwanzig-Caldeira-Leggett model provides the necessary framework for testing phenomenological schemes. It shows that averaging the conventional Hamiltonian of a free harmonic oscillator, $\tfrac{\langle P^{2}\rangle}{2M}+\tfrac{M}{2}\Omega^{2}\langle Q^{2}\rangle$, with the full equilibrium density matrix of the oscillator and its environment yields the correct expression for the energy of a damped oscillator when the damping function is temperature-independent [57, 63]. In this case, the resulting expression for the free energy remains valid even in the presence of temperature-dependent damping $\gamma(\omega,T)$.
For distance-dependent parameters, both the averaged potential and kinetic energies become distance dependent and jointly contribute to the interaction force. A similar statement is well known in the theory of van der Waals forces between atoms and molecules [84]. The phenomenological approach of Ref. [85] to the fluctuation-induced energy of the series RLC circuit is based precisely on the expression $\tfrac{1}{2}L\langle I^{2}\rangle+\tfrac{\langle Q_{C}^{2}\rangle}{2C}$, which includes both electric and magnetic contributions. The corresponding finite interaction force in the circuit was obtained in Refs. [86],  [^4] See the footnote on p.407 in Ref. [8]. and shown to be consistent with the general theory of Casimir and van der Waals forces. This force coincides with (7), (8) up to a minor typo  [^5] In the limit of large dissipation, described by equation (5.422) on page 248 in Ref. [86], the resistance $R$ under the logarithm sign must be squared. and therefore the theory developed in this manuscript supports the specific results obtained in Refs. [85],[86],[87].
An alternative phenomenological approach to Casimir-Lifshitz forces induced in electrical circuits was proposed in Ref. [89]. The principal difference from Refs. [85],[86],[87] is that Ref. [89] determines the forces solely from the potential term of the Hamiltonian, averaged over the state of the entire dissipative circuit, $\tfrac{\langle Q_{C}^{2}\rangle}{2C}$, neglecting all other Hamiltonian terms. This is inconsistent with the Zwanzig-Caldeira-Leggett model, insufficient for determining the forces, and leads to incorrect results which are in contradiction to those obtained here.
The discrepancy concerns both the explicit formulas for the forces in particular circuits, and the more general results and conclusions. One conclusion in Ref. [89] is that, in the dissipationless limit at zero temperature, the interaction potential differs from the zero-point energy. However, as follows from the damped oscillator theory in this limit, the interaction potential coincides with the zero-point energy when all relevant Hamiltonian terms are included. Another conclusion in Ref. [89] is that, within the lumped-element description, divergence of the interaction potential necessarily entails a divergence of the interaction force, even when it is only the resonant frequency that depends on distance. In particular, the expression for the interaction force in the parallel RLC circuit following from Ref. [89] diverges when $L$ is distance dependent at fixed $R$ and $C$. This contradicts our results (2) and (3) when applied to that case (see also (S24)-(S27) [65]). In the series RLC circuit, the averaged potential $\tfrac{\langle Q_{C}^{2}\rangle}{2C}$ is finite and leads to a finite contribution to the force [89], whereas the averaged magnetic energy $\tfrac{1}{2}L\langle I^{2}\rangle$, omitted in Ref. [89], diverges within the lumped circuit description. In agreement with Refs. [86],[87] and equations (7), (8) (see also (S20)-(S23) [65]), the corresponding generally divergent conribution to the force becomes finite when $C$ is distance-dependent at fixed $R$ and $L$.
V Conclusions